# HD-EPIC: A Highly-Detailed Egocentric Video Dataset

### v1

@@ Abstract
We show the potential of our highly-detailed annotations through a challenging VQA benchmark of 26K questions assessing the capability to recognise recipes, ingredients, nutrition, fine-grained actions, 3D perception, object motion, and gaze direction. The powerful long-context Gemini Pro only achieves 38.5% on this benchmark, showcasing its difficulty and highlighting shortcomings in current VLMs. We additionally assess action recognition, sound recognition, and long-term video-object segmentation on HD-EPIC.
@@ 4.3 Digital Twins: Scene & Object Movements
Object-Scene Interactions. With the 3D object locations, we associate locations with the closest fixture $f{\in}F$, subject to fixture-specific heuristics (e.g. objects must be within a counter’s x-y plane). We manually verify assignments and find the accuracy is 98%. On average objects move between 1.8 different fixtures per video (see Supp. C for stats).
@@ 4.4 HD-EPIC vs Prior Egocentric Datasets
Tab. 1 compares HD-EPIC to other egocentric datasets (full table in Supp.). Compared to the the largest dataset with labelled 3D environments (Aria Digital Twin [50]), HD-EPIC contains 5x more footage; has more annotations; and importantly was collected in an unscripted manner in the participants’ homes. In particular, HD-EPIC is the first to annotate recipes, nutritional values, detailed action segments, gaze and audio labels on the same set of videos. With these diverse and dense annotations, HD-EPIC constitutes a true zero-shot benchmark for video understanding.
@@ 5.1 HD-EPIC VQA Benchmark and Analysis
VQA Results Per Category and Per Prototype. Tab. 2 provides overall and per-category accuracy averaged over the prototype results shown in Fig. 11. Both language-only models only achieve 26.5% and 26.7%, only 6.7% above random. Open-source video VLMs (VideoLlama, LongVA, LlaVA-Video) perform similarly (27.4%, 29.3%, 32.4%) but have different strengths as shown in Fig. 11. For example, Llama better estimates nutrition, while the video is necessary to get above random performance on action recognition and gaze estimation. Gemini achieves the best performance, particularly for Recipe and Ingredient where external knowledge helps. However, the average performance (38.5%) and the gap to our sample human baseline (90.3%) shows the challenge posed by our VQA benchmark.
@@
Video Length. Fig. 12 shows models struggle with all video input lengths but are worst with 1-30 minute inputs.
@@ 5.2 Recognition Benchmarks
Sound Recognition. We evaluate 3 audio models [27, 34, 9], all trained on EPIC-Sounds. Tab. 4 shows a large gap in performance comparing HD-EPIC to EPIC-Sounds for SSAST (-28.4), ASF (-25.9) and TIM (-26.4). This shows audio is not as robust to new scenes as expected.
@@ 5.3 Long-Term VOS Benchmark
We construct a long-term video object segmentation benchmark using our segmentations and track associations (Sec. 9). Our benchmark has 400 sequences each with 1-3 objects and 2 hand masks (see Supp. D for details). While we have a lot more tracks, we keep it comparable to current benchmarks in size. We evaluate two models [63, 12] with a naive baseline where object masks are kept static. Fig. 14 shows the results. SAM2 [63] surpasses Cutie [12] for hands, but performs similarly poorly for objects, demonstrating the challenge of diversity in object perspective, lighting, location and occlusion.
@@ 6 Onwards…
Our dataset will be publicly available, with an open license – i.e. the videos, audio, gaze, camera pose estimates. All annotations will be published including Blender digital twin, detailed narrations, object movements, object masks and 3D locations, long-object tracks, and object-action-fixture labels. All four benchmarks will be published with starter code. We hope HD-EPIC will direct future research to a more holistic perception of egocentric videos.
@@ Acknowledgements
The project was supported by a unrestricted charitable donation from Meta (Aria Project Partnership) to the University of Bristol. Gemini Pro results are supported by a research credits grant from Google DeepMind. Research at Bristol is supported by EPSRC Fellowship UMPIRE (EP/T004991/1), EPSRC Program Grant Visual AI (EP/T028572/1) and EPSRC Doctoral Training Program. O Emara, K Flanagan and F Abdelazim are supported by UKRI CD in Interactive AI (EP/S022937/1). Research at Leiden is supported by the Dutch Research Council (NWO) under a Veni grant (VI.Veni.222.160). Research at Singapore is supported by Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (No. MSS23C018).
We thank Rajan from Elancer and his team, for their huge assistance with temporal and audio annotation. We thank Srdjan Delic and his team for their assistance with mask annotations and object associations. We also thank Owen Tyley for the 3D Digital Twin of the kitchen environments using Blender.
We thank David Fouhey and Evangelos Kazakos for early feedback on the project. We thank Pierre Moulon, Vijay Baiyya and Cheng Peng from the Aria team for technical assistance in using the MPS code and services. We acknowledge the usage of Ismbard AI Phase 1 and EPSRC Tier-2 Jade clusters.
@@ B.4 Post-Processing: Multi-Video SLAM and Gaze
Dataset
Val&Test
Hours
Action
segments
Unscripted
Free
setting
Recipe Nutrition Gaze
Audio
labels
Object
Hand
3D object
over time
Labelled 3D
environment
Camera
pose
Fully
annotated
@@ C.3 Digital Twin
Hand and Object Segmentations To improve performance, we first manually annotate a small set of images from each video (5-10 frames) for both left and right hands. To cover the diverse locations and lighting conditions, we use the camera pose estimates to identify clusters of hotspots and select a set of images that represent hands in action at various locations. This is sent to dedicated annotators who manually segment these. We follow the same approach in [19], where hands and arms are segmented jointly to avoid the ad hoc decision of the hand boundary. We use the TORAS [33] annotation tool for all our manual segmentation. For the memory, we manually annotate 768 frames across 156 videos, which accounts for roughly $0.02\%$ of the total frames.
@@ D.1.6 Object Motion
Stationary Object Localization. Here, we evaluate the capability to perform the reverse task: determining when an object remains static in the video. The question is: ”After the object <BBOX Y1 X1 Y2 X2> seen at <TIME HH:MM:SS video 1> is first moved, from which of the following starting times does the object remain static for more than <X> seconds?”. Similar to the previous task, the input is the entire video, and the object is specified by its bounding box at a given timestamp. The static duration is determined using long-term object tracking annotations (Sec. C.3). We find $t_{1}$ and $t_{2}$, the two longest static durations of an object. A value is then randomly sampled from the range $(t_{2}+5,t_{1})$ to fill the placeholder <X>, which specifies the required static duration for the question. Objects with a longest static durations of less than 5 seconds and less than three movements are excluded. We also use the criteria mentioned in the previous question to filter objects with rapid movements and spatial inconsistency across tasks. For the positive answer, a timestamp is randomly selected from the interval starting at the beginning of the maximum static duration and ending at a point determined by subtracting the previously chosen static duration from its endpoint. Negative answers are randomly sampled from the video from the start of the video to the last movement of the object, ensuring no overlap with the positive. All answers are provided as timestamps formatted as <TIME HH:MM:SS video 1>.
@@ D.2 Additional VQA Experiments
Making predictions from the full ground-truth (GT) narrations is better than using GT action clusters for all categories demonstrating the benefit of more detailed short-term understanding. Using GT narrations instead of visual input gives better performance for all question categories, highlighting the challenge of visual understanding in HD-EPIC. Having GT narrations helps most with recipe, ingredient and fine-grained action questions but far from solves these categories e.g. ingredient is 48.6%. Regardless, nutrition, 3D perception, object motion, and gaze still have low performance with GT narrations demonstrating their difficulty.
@@ D.4 Long-Term VOS Benchmark
DAVIS [53] 30 2.8s 83.3s 2.03 1,999
YouTube-VOS [82] 474 4.3s 2,037.3s 1.89 12,593
Ours 400 533.9s 38,785.7s 4.16 6,692
@@
The details of our long-term VOS benchmark are compared against the validation sets of two popular VOS benchmarks, DAVIS [53] and YouTube-VOS [82]. Our benchmark includes 400 sequences, with an average duration of 533.9 seconds. This is calculated using the first and last frame indices of the annotated frames. It is clear that our benchmark provides a much longer temporal duration for video object segmentation evaluation. Although the number of annotated frames remains lower than YouTube-VOS, there is scope to perform VOS on a much longer set of object movements. Furthermore, the table shows that the average number of objects in each sequence is more than doubled compared to the prior benchmarks.

### v2

@@ Abstract
We show the potential of our highly-detailed annotations through a challenging VQA benchmark of 26K questions assessing the capability to recognise recipes, ingredients, nutrition, fine-grained actions, 3D perception, object motion, and gaze direction. The powerful long-context Gemini Pro only achieves 37.6% on this benchmark, showcasing its difficulty and highlighting shortcomings in current VLMs. We additionally assess action recognition, sound recognition, and long-term video-object segmentation on HD-EPIC.
@@ 4.3 Digital Twins: Scene & Object Movements
Object-Scene Interactions. With the 3D object locations, we associate locations with the closest fixture $f{\in}F$, subject to fixture-specific heuristics (e.g. objects must be within a counter’s x-y plane). We manually verify all assignments, correcting any errors. On average objects move between 1.8 different fixtures per video (see Supp. C for stats).
@@ 4.4 HD-EPIC vs Prior Egocentric Datasets
Tab. 1 compares HD-EPIC to other egocentric datasets (full table in Supp.). Compared to the largest dataset with labelled 3D environments (Aria Digital Twin [50]), HD-EPIC contains 5x more footage; has more annotations; and importantly was collected in an unscripted manner in the participants’ homes. In particular, HD-EPIC is the first to annotate recipes, nutritional values, detailed action segments, gaze and audio labels on the same set of videos. With these diverse and dense annotations, HD-EPIC constitutes a true zero-shot benchmark for video understanding.
@@ 5.1 HD-EPIC VQA Benchmark and Analysis
VQA Results Per Category and Per Prototype. Tab. 2 provides overall and per-category accuracy averaged over the prototype results shown in Fig. 11. Both language-only models only achieve 26.5% and 26.7%, only 6.7% above random. Open-source video VLMs (VideoLlama, LongVA, LlaVA-Video) perform similarly (27.4%, 29.3%, 32.4%) but have different strengths as shown in Fig. 11. For example, Llama better estimates nutrition, while the video is necessary to get above random performance on action recognition and gaze estimation. Gemini achieves the best performance, particularly for Recipe and Ingredient where external knowledge helps. However, the average performance (37.6%) and the gap to our sample human baseline (90.3%) shows the challenge posed by our VQA benchmark.
@@
Video Length. Fig. 12 shows models struggle with all video lengths but are worst with inputs $\geq$1 minute.
@@ 5.2 Recognition Benchmarks
Sound Recognition. We evaluate 3 audio models [27, 34, 9], all trained on EPIC-Sounds. Tab. 4 shows a large gap in performance comparing HD-EPIC to EPIC-Sounds for SSAST (-28.4), ASF (-25.9) and TIM (-26.4). This shows audio is not sufficiently robust to new scenes or devices.
@@ 5.3 Long-Term VOS Benchmark
We construct a long-term video object segmentation benchmark using our segmentations and track associations (Sec. 9). Our benchmark has 1000 sequences, each with 1-5 objects and 2 hand masks (see Supp. D for details). While we have a lot more tracks, we keep it comparable to current benchmarks in size. We evaluate two models [64, 12] with a naive baseline where object masks are kept static. Fig. 14 shows the results. SAM2 [64] surpasses Cutie [12] for hands, but does worse on objects. Overall, objects have added challenge in diversity in perspective, lighting, location and occlusion.
@@ 6 Onwards…
HD-EPIC is available from: http://dx.doi.org/10.5523/bris.3cqb5b81wk2dc2379fx1mrxh47i.e. the videos, audio, gaze, blender digital twin, camera pose estimates. Annotations are available at: http://hd-epic.github.ioi.e. object movements, object masks and 3D locations, long-object tracks, and object-action-fixture labels. We hope HD-EPIC will direct future research to a more holistic perception of egocentric videos.
@@ Acknowledgements
Research at Bristol is supported by EPSRC Fellowship UMPIRE (EP/T004991/1), EPSRC Program Grant Visual AI (EP/T028572/1) and EPSRC Doctoral Training Program. O Emara, K Flanagan and F Abdelazim are supported by UKRI CD in Interactive AI (EP/S022937/1). The project is also supported by a unrestricted charitable donation from Meta (Aria Project Partnership) to the University of Bristol. Gemini Pro results are supported by a research credits grant from Google DeepMind.
Research at Leiden is supported by the Dutch Research Council (NWO) under a Veni grant (VI.Veni.222.160). Research at Singapore is supported by Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (No. MSS23C018).
We thank Rajan from Elancer and his team, for their huge assistance with temporal and audio annotation. We thank Srdjan Delic and his team for their assistance with mask annotations. We thank Damian Steer and the University of Bristol’s RDSF team for hosting and maintaining the dataset. We also thank Owen Tyley for the 3D Digital Twin of the kitchen environments using Blender.
We thank David Fouhey and Evangelos Kazakos for early feedback on the project. We thank Pierre Moulon, Vijay Baiyya and Cheng Peng from the Aria team for technical assistance in using the MPS code and services.
We acknowledge the usage of EPSRC Tier-2 Jade clusters. The authors also acknowledge the use of Isambard-AI National AI Research Resource (AIRR). Isambard-AI is operated by the University of Bristol and is funded by the UK Government’s Department for Science, Innovation and Technology (DSIT) via UK Research and Innovation; and the Science and Technology Facilities Council [ST/AIRR/I-A-I/1023].
@@ B.4 Post-Processing: Multi-Video SLAM and Gaze
Hours Segments Setting Labels over time environment pose annotated
@@ C.3 Digital Twin
Hand and Object Segmentations To improve performance, we first manually annotate a small set of images from each video (5-10 frames) for both left and right hands. To cover the diverse locations and lighting conditions, we use the camera pose estimates to identify clusters of hotspots and select a set of images that represent hands in action at various locations. This is sent to dedicated annotators who manually segment these. We follow the same approach in [19], where hands and arms are segmented jointly to avoid the ad hoc decision of the hand boundary. We use the TORAS [33] annotation tool for all our manual segmentation. We manually annotate 768 frames across 156 videos, which accounts for roughly $0.02\%$ of the total frames.
@@ D.1.6 Object Motion
Stationary Object Localization. Here, we evaluate the capability to perform the reverse task: determining when an object remains static in the video. The question is: ”After the object <BBOX Y1 X1 Y2 X2> seen at <TIME HH:MM:SS video 1> is first moved, from which of the following starting times does the object remain static for more than <X> seconds?”. Similar to the previous task, the input is the entire video, and the object is specified by its bounding box at a given timestamp. The static duration is determined using long-term object tracking annotations (Sec. C.3). We find $t_{1}$ and $t_{2}$, the two longest static durations of an object. A value is then randomly sampled from the range $(t_{2}+5,t_{1})$ to fill the placeholder <X>, which specifies the required static duration for the question. We also use the criteria mentioned in the previous question to filter objects with rapid movements and spatial inconsistency across tasks. For the positive answer, a timestamp is randomly selected from the interval starting at the beginning of the maximum static duration and ending at a point determined by subtracting the previously chosen static duration from its endpoint. Negative answers are randomly sampled from the video from the start of the video to the last movement of the object, ensuring no overlap with the positive. All answers are provided as timestamps formatted as <TIME HH:MM:SS video 1>.
@@ D.2 Additional VQA Experiments
Making predictions from the full ground-truth (GT) narrations is better than using GT action clusters for all categories demonstrating the benefit of more detailed short-term understanding. Using GT narrations instead of visual input gives better performance for all question categories, highlighting the challenge of visual understanding in HD-EPIC. Having GT narrations helps most with recipe and fine-grained action and object motion questions but far from solves these categories e.g. fine-grained action is 63.6%. Regardless, nutrition, 3D perception, object motion, and gaze still have low performance with GT narrations demonstrating their difficulty.
@@ D.4 Long-Term VOS Benchmark
DAVIS [57] 30 2.8s 83.3s 2.0 1,999
YouTube-VOS [84] 507 21.9s 11,097.0s 2.1 13,710
Ours 1000 561.0s 561,034.1s 6.3 20,548
@@
The details of our long-term VOS benchmark are compared against the validation sets of two popular VOS benchmarks, DAVIS 2017 [57] and YouTube-VOS 2019 [84]. Our benchmark includes 1000 sequences, with an average duration of 561.0 seconds. This is calculated using the first and last frame indices of the annotated frames. When calculating durations for the DAVIS and YouTube-VOS we assume FPS values of 24 and 6, respectively, as stated in the original papers [53, 83]. It is clear that our benchmark provides a much longer temporal duration for video object segmentation evaluation. Furthermore, the table shows that the average number of objects in each sequence is nearly tripled compared to the prior benchmarks.